Find the slope-intercept equatin of the line that passes through the point (6,4) and has slope -2
Answer by mysterynorth(5)

I don't know what your solving for but here's what I have fr you.
plug in the points
4(which is really y) = -2 (slope) (6) +b
4= -2 (6) +b
multiply slope times x ( -2 * 6)
Add 12 to each side (which is your answer for -2* 6's opposite so they cancle eachother out)
your left with 16= B go back and plug b in
your left with either
y = -2 (x) + 16 or
4 = -2 (6) + 16